OpenStudy (anonymous):

whats the solve for 3u+4=-17

OpenStudy (anonymous):

solve for u?

OpenStudy (anonymous):

3u=-17-4 1/3(3u=-21)1/3 u=-7

OpenStudy (anonymous):

thanks <3